# PseudoNeg-MAE: Self-Supervised Point Cloud Learning using Conditional Pseudo-Negative Embeddings

### v1

@@ III-D COPE Network Architecture
First, we embed $g$ into a harmonic space $H(\cdot)$, using a set of high-frequency functions [25]. Subsequently, these embedded quaternions are processed through an MLP network to extract transformation-sensitive embeddings, denoted as $\mathbf{h}_{g}^{\prime}\in\mathbb{R}^{\frac{d}{k}}$. Independently estimating each element of $\Theta_{g}$ is computationally expensive. Therefore, we utilize shared learnable embedding vectors $\Psi\in\mathbb{R}^{\frac{d}{k}\times d}$, which are responsible for the entries of the rows in $\Theta_{g}$. We perform an element-wise multiplication of these vectors with our $h_{g}^{\prime}$, i.e., $\Gamma=h_{g}^{\prime}\odot\Psi$, where $\Gamma\in\mathbb{R}^{d\times\frac{d}{k}}$. A shared non-linear MLP network is then used to expand each column of $\Gamma$ to the dimension $d$. This process yields the linear weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. Consequently, $\Theta_{g}$ is determined solely based on the input transformation.
@@
We generate pseudo negatives using COPE by sampling a set of random rotations $\{{g_{r}}\}_{r=1}^{M}$. We uniformly sample quaternions representing 3D rotations to ensure coverage over the rotation space. Specifically, for each pseudo negative, we generate a random vector $\mathbf{v}_{r}=(w_{r},x_{r},y_{r},z_{r})$ by sampling from a standard normal distribution $\mathbf{v}_{r}\sim\mathcal{N}(0,I_{4})$. We then normalize $\mathbf{v}_{r}$ to obtain a unit quaternion $g_{r}$ and ensure the scalar part is nonnegative:
$g_{r}=\frac{\mathbf{v}_{r}}{|\mathbf{v}_{r}|}\cdot\text{sign}(w_{r}).$ (9)
This approach ensures that the set of quaternions ${g_{r}}$ is uniformly distributed over the space of 3D rotations. Random rotations are fed to the COPE for generating pseudo negatives as in Eq. 6.
III-E PseudoNeg-MAE Transformer Backbone
@@
Our point cloud transformer backbone is based on a Siamese encoder-decoder network, denoted as $f$, in which there are two branches with shared weights. PseudoNeg-MAE operates on pairs of point clouds: a randomly sampled point cloud $\mathbf{x}_{i}$ and its transformed point cloud $\mathbf{x}_{i}^{+}$ with the applied transformation $g$. Following the standard transformer architecture on the point cloud, we first extract a sequence of non-overlapping point patches, which are then converted to a set of tokens. Given an input point cloud $\mathbf{x}$, we first used Furthest Point Sampling (FPS) to sample a set of $n$ center points. Then, K Nearest-Neighbour (KNN) is utilized to select $k$ neighboring points around each center point to construct $n$ point patches. Following this, patch tokens are generated using a mini-PointNet [3] and fed to the transformer encoder. However, in PseudoNeg-MAE, not all the patch tokens are processed by the encoder. A random masking strategy is applied independently to $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, masking $m$ number of the point patches. Only visible tokens are fed to an encoder ($\mathcal{E}$). This independent masking strategy encourages the learning of representations that are robust to noise and occlusions.
Then, the encoded embeddings of visible tokens from the two inputs are sent to the decoders with shared weights. An extra learnable [CLS] token is appended to each encoder output. We implement a lightweight decoder $\mathcal{D}$. The objective of the decoder is to aggregate the encoder outputs into representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$.
Figure 2: Conceptual visualization of the loss $\mathcal{L}_{\textit{PseudoNeg-MAE}{}}$. Using COPE, we generate the anchor $\Theta_{g}\mathbf{z}_{i}$ and a set of pseudo negatives $\Theta_{g_{r}}\mathbf{z}_{i}$ for the correspondoing positive $\mathbf{z}_{i}^{+}$. As per our main goal, due to these pseudo negatives, the model remains sensitive to different transformations rather than becoming completely invariant to them.
III-F Relative Pose Estimation Using COPE
@@
To tackle the challenge of relative pose estimation between source and target point clouds, we introduce a novel algorithm that utilizes the COPE network to optimize the transformation parameters directly. The algorithm focuses on estimating the quaternions to minimize the difference between the transformed source and target embeddings, thus achieving precise alignment.
@@
The process, detailed in Algorithm 1, starts by exploring a set of randomly generated quaternions $\{g_{r}\}$, representing potential rotations between the source and target. Each quaternion is evaluated in a bi-directional manner—applying the transformation to both the source and the target embeddings via $g_{r}$ and $g_{r}^{-1}$. Here $g_{r}^{-1}$ is the inverse of $g_{r}$. The quaternions are refined through an iterative process, where each is adjusted according to the gradient update. After updating the quaternions, it is essential to ensure they remain valid unit quaternions, which is achieved by the normalization step at line 1. The quaternion associated with the minimum loss is designated as $g_{est}$ serving as the final solution for the alignment. This process ensures that the iterative refinement systematically approaches the global minimum of the loss landscape. In summary, this algorithm extends the capabilities of COPE to solve relative pose estimation by optimizing quaternions to align global embeddings of point clouds.
Data: Learning rate $\epsilon=0.01$
Data: Source and target embeddings $\mathbf{z}_{src},\mathbf{z}_{tgt}$
1 Initialize $\mathcal{L}_{min}=\infty$
2for j in $N_{out}$ do
3       Initialize $g_{r}=(w_{r},x_{r},y_{r},z_{r})$
4       for k in $N_{in}$ do
5             $\mathbf{z}^{g_{r}}=\mathrm{\text{COPE}{}}(g_{r})\mathbf{z}_{src}$
6             $\mathbf{z}^{g_{r}^{-1}}=\mathrm{\text{COPE}{}}(g_{r}^{-1})\mathbf{z}_{tgt}$
7             $\mathcal{L}_{g_{r}}=\left\|\mathbf{z}_{tgt}-\mathbf{z}^{g_{r}}\right\|_{2}^{2}+\left\|\mathbf{z}_{src}-\mathbf{z}^{g_{r}^{-1}}\right\|_{2}^{2}$
8             $g_{r}^{*}=g_{r}-\epsilon\nabla_{g_{r}}L_{g_{r}}$
9             $g_{r}=\frac{g_{r}}{\|g_{r}\|}\cdot\text{sign}(w_{r}^{*})$
@@ 8             $g_{r}^{*}=g_{r}-\epsilon\nabla_{g_{r}}L_{g_{r}}$
11
12      if $\mathcal{L}_{g_{r}}<\mathcal{L}_{min}$ then
13             Update $\mathcal{L}_{min}=\mathcal{L}_{g_{r}}$, set $g_{est}=g_{r}$
@@
return $g_{est}$
Algorithm 1 Optimization of Transformation Estimation with Minimum Loss Tracking.
IV Experiments
IV-A Pre-training and implementation details
The encoder $\mathcal{E}$ comprises 12 layers, each with an internal dimension of 384. The decoder $\mathcal{D}$ is implemented as a single transformer layer with [CLS] token. Training is conducted over 1600 epochs with a batch size of 512. We utilize the AdamW optimizer, starting with a learning rate of $1\times 10^{-4}$ and applying a cosine annealing with a linear warmup scheduler. After thorough validation of the ShapeNet dataset, we set the $\beta$ parameter of the loss function to 0.3 and the number of pseudo negatives to eight.
TABLE I: Shape classification results on ModelNet40 dataset and ScanObjNN’s OBJ-BG subset under z/z and S0(3)/SO(3) evaluation scenario.
Method ModelNet40 ScanObjNN–OBJ-BG
@@ IV-A Pre-training and implementation details
TABLE II: Relative pose estimation results over five categories with E2PN protocol. Here all the other compared methods are supervised. Among them, KPConv is a non-equivariant method while EPN and E2PN are equivariant networks. Our self-supervised approach performs better than the supervised equivariant methods.
@@
V Results on Downstream Tasks
We evaluate our model on downstream tasks, specifically shape classification and relative pose estimation for both objects and scene levels, to demonstrate the advantage of discriminative and rotation-sensitive properties of our approach.
V-A Shape Classification
We evaluate our model on two prominent 3D object classification datasets, namely, ModelNet40 and ScanObjectNN. ModelNet40 [42] consists of 12,311 clean 3D CAD models from 40 categories, split into 9,843 instances for training and 2,468 for testing. ScanObjectNN [43] is a more challenging dataset derived from real-world indoor scenes containing approximately 15,000 objects across 15 categories. It includes three variants: OBJ-ONLY with isolated segmented objects, OBJ-BG featuring objects with contextual background elements, and OBJ-T50-RS with perturbations in object extraction.
We follow the protocols of [44] for the evaluation in Tab. I: z/z – training and testing under rotations around the z-axis and SO(3)/SO(3) – training and testing the under the arbitrary 3D rotations. Tab. I provides a detailed comparison of different point cloud backbones applied to the ModelNet40 and ScanObjectNN datasets across different protocols. The table is divided into sections detailing Supervised and SSL approaches. Here supervised section consists of equivariant methods [31] - [38] which are designed to handle rotations. PseudoNeg-MAE, our SSL-based model surpasses the performance of both supervised and self-supervised methods. Specifically, in the challenging SanObjNN OBJ-BG scenario, PseudoNeg-MAE achieves 89.0%, which is a significant improvement over the best supervised method and leads other SSL methods by a considerable margin. This demonstrates that our novel pretraining approach enhances the generalization capabilities across diverse ranges.
V-B Relative Pose Estimation
We assess our model on relative pose estimation benchmark tasks, where the goal is to predict the relative rotation between pairs of point clouds. We fine-tune the network $f$ and COPE using $\mathcal{L}_{\text{align}^{*}}$ and $\mathcal{L}_{\text{cope}}$, and conduct inference following Algorithm 1. The evaluation is conducted under two scenarios: object-level and scene-level point clouds. We use mean isotropic rotation error [45] for the evaluation.
Object Level
In the object-level evaluation, we utilize five categories from the ModelNet40 dataset, following the protocol established by EPN [40]. We present both quantitative and qualitative evaluations in Tab. II and Fig. 3, respectively. Our self-supervised model, PseudoNeg-MAE, shows the best overall performance compared to supervised counterparts, including both non-equivariant (such as KPConv [39]) and equivariant (EPN [40] and E2PN  [41]) networks. By achieving lower maximum rotation errors across all categories, PseudoNeg-MAE demonstrates its superior ability to learn robust rotation-sensitive embeddings. This helps to significantly enhance the alignment accuracy of point clouds, validating the effectiveness of our proposed method in handling all ranges of rotations.
Figure 3: Illustration of relative pose estimation between source (yellow) and target (blue) point clouds with for KPConv, EPN, E2PN, and PseudoNeg-MAE. Red indicates transformed source clouds.
Scene Level
@@ V-B Relative Pose Estimation
In addition to the object-level evaluation, we extend our analysis to relative pose estimation at the scene level using the 7Scenes dataset, a real-world indoor dataset that is a subset of 3DMatch [48]. This dataset features over 100,000 points per scene. We sample 1024 points to form the source and target point clouds. We compare our method with correspondence-free pose estimation methods where FMR [46] is an iterative approach and EquivReg [47] uses an equivariant encoder. Following the protocol described EquivReg [47], we generate random rotations by uniformly sampling the angles with the specified maximum $\left(0^{\circ},\ldots,180^{\circ}\right)$, alongside randomly selected rotation axes, to compute rotation errors. The findings are summarized in Fig. 4 with a logarithmic scale for better visualization. It shows the PseudoNeg-MAE’s ability to maintain consistent lower error across the range of angles.
VI Conclusion

### v2

@@ 3 Proposed Approach - PseudoNeg-MAE
COPE Network Architecture: COPE generates linear weights conditioned on the transformation set ${\{g}\}$. In our work, we consider rotations as transformations and utilize quaternions to represent them as inputs to the COPE. The network architecture is illustrated in the right section of Fig. 1. First, we embed $g$ into a harmonic space $H(\cdot)$, using a set of high-frequency functions [30]. Subsequently, these embedded quaternions are processed through an MLP network to extract transformation-sensitive embeddings, denoted as $\mathbf{h}_{g}^{\prime}\in\mathbb{R}^{\frac{d}{k}}$. Independently estimating each element of $\Theta_{g}$ is computationally expensive. Therefore, we utilize shared learnable embedding vectors $\Psi\in\mathbb{R}^{\frac{d}{k}\times d}$, which are responsible for the entries of the rows in $\Theta_{g}$. We perform an element-wise multiplication of these vectors with our $h_{g}^{\prime}$, i.e., $\Gamma=h_{g}^{\prime}\odot\Psi$, where $\Gamma\in\mathbb{R}^{d\times\frac{d}{k}}$. A shared non-linear MLP network is then used to expand each column of $\Gamma$ to the dimension $d$. This process yields the linear weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. Consequently, $\Theta_{g}$ is determined solely based on the input transformation.
@@
We generate pseudo negatives using COPE by sampling a set of random rotations $\{{g_{r}}\}_{r=1}^{M}$. We uniformly sample quaternions representing 3D rotations to ensure coverage over the rotation space. Specifically, for each pseudo negative, we generate a random vector $\mathbf{v}_{r}=(w_{r},x_{r},y_{r},z_{r})$ by sampling from a standard normal distribution $\mathbf{v}_{r}\sim\mathcal{N}(0,I_{4})$. We then normalize $\mathbf{v}_{r}$ to obtain a unit quaternion $g_{r}$ and ensure the scalar part is nonnegative: $g_{r}=\frac{\mathbf{v}_{r}}{|\mathbf{v}_{r}|}\cdot\text{sign}(w_{r}).$ This approach ensures that the set of quaternions ${g_{r}}$ is uniformly distributed over the space of 3D rotations. Random rotations are fed to the COPE for generating pseudo negatives as in Eq. 2.
PseudoNeg-MAE Transformer Backbone: Our contrastive framework is built on the Masked AutoEncoder (MAE), combining the strengths of both predictive (MAE) and discriminative (contrastive) model training paradigms. We employ a shared Siamese encoder with a point cloud transformer backbone $f$, initialized with weights from an MAE in which there are two branches with shared weights. PseudoNeg-MAE operates on pairs of point clouds: a randomly sampled point cloud $\mathbf{x}_{i}$ and its transformed point cloud $\mathbf{x}_{i}^{+}$ with the applied transformation $g$. Following the standard transformer architecture on the point cloud, we first extract a sequence of non-overlapping point patches, which are then converted to a set of tokens.
@@
Given an input point cloud $\mathbf{x}$, we first used Furthest Point Sampling (FPS) to sample a set of $n$ center points. Then, K Nearest-Neighbour (KNN) is utilized to select $k$ neighboring points around each center point to construct $n$ point patches. Following this, patch tokens are generated using a mini-PointNet [3] and fed to the transformer encoder. However, in PseudoNeg-MAE, not all the patch tokens are processed by the encoder. A random masking strategy is applied independently to $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, masking $m$ number of the point patches. Only visible tokens are fed to an encoder ($\mathcal{E}$). This independent masking strategy encourages the learning of representations that are robust to noise and occlusions. Then, the encoded embeddings of visible tokens from the two inputs are sent to the decoders with shared weights. An extra learnable [CLS] token is appended to each encoder output. We implement a lightweight decoder $\mathcal{D}$. The objective of the decoder is to aggregate the encoder outputs into representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$.
@@
Relative Pose Estimation Using COPE: To tackle the challenge of relative pose estimation between source and target point clouds, we introduce a novel algorithm that utilizes the COPE network to optimize the transformation parameters directly. The algorithm focuses on estimating the quaternions to minimize the difference between the transformed source and target embeddings, thus achieving precise alignment.
@@
The process, detailed in Algorithm 1, starts by exploring a set of randomly generated quaternions $\{g_{r}\}$, representing potential rotations between the source and target. Each quaternion is evaluated in a bi-directional manner—applying the transformation to both the source and the target embeddings via $g_{r}$ and $g_{r}^{-1}$. Here $g_{r}^{-1}$ is the inverse of $g_{r}$. The quaternions are refined through an iterative process, where each is adjusted according to the gradient update weighted by $\epsilon=0.01$. After updating the quaternions, it is essential to ensure they remain valid unit quaternions, which is achieved by the normalization step at line 1. The quaternion associated with the minimum loss is designated as $g_{est}$ serving as the final solution for the alignment. This process ensures that the iterative refinement systematically approaches the global minimum of the loss landscape. In summary, this algorithm extends the capabilities of COPE to solve relative pose estimation by optimizing quaternions to align global embeddings of point clouds.
Data: Source and target $\mathbf{z}_{src},\mathbf{z}_{tgt}$
1 Initialize $\mathcal{L}_{\min}=\infty$;
2 for $j\!=\!1\ldots N_{\text{out}}$ do
3       Initialise $g_{r}=(w_{r},x_{r},y_{r},z_{r})$;
4       for $k\!=\!1\ldots N_{\text{in}}$ do
5             $\mathbf{z}^{g_{r}}\!=\!\frac{\text{COPE}(g_{r})\mathbf{z}_{src}}{\|\text{COPE}(g_{r})\mathbf{z}_{src}\|}\;$;
6             $\mathbf{z}^{g_{r}^{-1}}\!=\!\frac{\text{COPE}(g_{r}^{-1})\mathbf{z}_{tgt}}{\|\text{COPE}(g_{r}^{-1})\mathbf{z}_{tgt}\|}\;$;
7             $\mathcal{L}_{g_{r}}\!=\!\|\mathbf{z}_{tgt}-\mathbf{z}^{g_{r}}\|_{2}^{2}+\|\mathbf{z}_{src}-\mathbf{z}^{g_{r}^{-1}}\|_{2}^{2}$;
8             $g_{r}\leftarrow g_{r}-\epsilon\nabla_{g_{r}}\mathcal{L}_{g_{r}}$;
9             $g_{r}\leftarrow\displaystyle\frac{g_{r}}{\|g_{r}\|}\,\text{sign}(w_{r})$;
10
@@ 8             $g_{r}\leftarrow g_{r}-\epsilon\nabla_{g_{r}}\mathcal{L}_{g_{r}}$;
12      if $\mathcal{L}_{g_{r}}<\mathcal{L}_{\min}$ then
13             $\mathcal{L}_{\min}\leftarrow\mathcal{L}_{g_{r}}$,$g_{\text{est}}\leftarrow g_{r}$;
@@
return $g_{\text{est}}$
Algorithm 1 Transformation Estimation with COPE.
Table 1: Shape classification results on ModelNet40 dataset and ScanObjNN’s OBJ-BG subset under z/z and S0(3)/SO(3) evaluation scenario.
Method ModelNet40 ScanObjNN
@@
4 Results
We evaluate our model on downstream tasks, specifically shape classification and relative pose estimation for both objects and scene levels, to demonstrate the advantage of the invariant and equivariant properties of our approach. Details of our hyperparameters and pre-training settings are given in the Supplementary Materials.
Shape Classification: We evaluate our model on two prominent 3D object classification datasets, namely, ModelNet40 and ScanObjectNN. ModelNet40 [43] consists of 12,311 clean 3D CAD models from 40 categories, split into 9,843 instances for training and 2,468 for testing. ScanObjectNN [44] is a more challenging dataset derived from real-world indoor scenes containing approximately 15,000 objects across 15 categories.
We follow the protocols of [45] for the evaluation in Tab. 3: z/z – training and testing under rotations around the z-axis and SO(3)/SO(3) – training and testing the under the arbitrary 3D rotations. Tab. 3 provides a detailed comparison of different point cloud backbones applied to the ModelNet40 and ScanObjectNN datasets across different protocols. The table is divided into sections detailing Supervised and SSL approaches. Here supervised section consists of equivariant methods [33] - [42] which are designed to handle rotations. Our SSL-based model, PseudoNeg-MAE, achieves superior performance on ModelNet40 and ScanObjNN, outperforming all supervised and self-supervised methods under S0(3)/SO(3) on challenging ScanObjNN OBJ-BG by a considerable margin, and ModelNet40 under z/z while remaining highly competitive under S0(3)/SO(3). This demonstrates that our novel pretraining approach enhances the generalization capabilities across diverse ranges.
Relative Pose Estimation: We assess our model on relative pose estimation benchmark tasks, where the goal is to predict the relative rotation between pairs of point clouds. We fine-tune the network $f$ and COPE using $\mathcal{L}_{\text{align}^{*}}$ and $\mathcal{L}_{\text{cope}}$, and conduct inference following Algorithm 1. The evaluation is conducted under two scenarios: object-level and scene-level point clouds. We use the mean isotropic rotation error [46] for the evaluation.
Table 2: Relative pose estimation results over five categories with E2PN protocol. Here, all the other compared methods are supervised. Among them, KPConv is a non-equivariant method while EPN and E2PN are equivariant networks. Our self-supervised approach performs better than the supervised equivariant methods.
@@ 4 Results
Object Level: In the object-level evaluation, we utilize five categories from the ModelNet40 dataset, following the protocol established by EPN [15]. We present both quantitative and qualitative evaluations in Tab. 2 and Fig. 3, respectively. Our self-supervised model, PseudoNeg-MAE, shows the best overall performance compared to supervised counterparts, including both non-equivariant (such as KPConv [17]) and equivariant (EPN [15] and E2PN  [16]) networks. By achieving lower maximum rotation errors across all categories, PseudoNeg-MAE demonstrates its superior ability to learn robust rotation-sensitive embeddings. This helps to significantly enhance the alignment accuracy of point clouds, validating the effectiveness of our proposed method in handling all ranges of rotations.
Figure 2: Illustration of relative pose estimation between source (yellow) and target (blue) point clouds for KPConv, EPN, E2PN, and PseudoNeg-MAE. Red indicates transformed source clouds.
@@
Scene Level: In addition to the object-level evaluation, we extend our analysis to relative pose estimation at the scene level using the 7Scenes dataset, a real-world indoor dataset that is a subset of 3DMatch [48]. This dataset features over 100,000 points per scene. We sample 1024 points to form the source and target point clouds. We compare our method with correspondence-free pose estimation methods where FMR [47] is an iterative approach and EquivReg [19] uses an equivariant encoder. Following the protocol described EquivReg [19], we generate random rotations by uniformly sampling the angles with the specified maximum $\left(0^{\circ},\ldots,180^{\circ}\right)$, alongside randomly selected rotation axes, to compute rotation errors. The findings are summarized in Fig. 3 with a logarithmic scale for better visualization. It shows the PseudoNeg-MAE’s ability to maintain consistent lower error across the range of angles.
5 Conclusion